OpenStudy (bbypanda16):

There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. (4 points)

OpenStudy (anonymous):

Ok so first you need to make functions to model these... Neighborhood A is an exponential function and B is a linear function Do you think you can do that?

OpenStudy (gnorris):

for part A neighborhood A f(x)=30+6x neighborhood B f(x)=45+3x where x = number of years

OpenStudy (bbypanda16):

Thank you <3